### v1

# Agent Contracts: A Formal Framework for Resource-Bounded Autonomous AI Systems (Full)

Abstract
[…]
1 Introduction
In late 2025, an engineering team deployed a multi-agent research system with four specialized agents. Two agents fell into a recursive clarification loop, running undetected for eleven days. When the invoice arrived, the team discovered a \$47,000 API bill [26]. The system had no stop conditions, no budget limits, and no real-time cost monitoring. This incident encapsulates a fundamental problem—we have built AI agents capable of autonomous action but lack formal mechanisms to bound their behavior.
Such failures reflect systemic gaps, not implementation bugs [10]. Gartner predicts that over 40% of agentic AI projects will be canceled by 2027 due to escalating costs or inadequate risk controls [15], even as agentic AI is projected to appear in 33% of enterprise software by 2028 [14]. A recent MIT Sloan study finds that 35% of organizations already deploy agentic AI, with leaders citing the tension between supervision and autonomy as a core challenge requiring centralized governance infrastructure [40]. Agents are becoming capable of sustained autonomous operation spanning hours or days [42, 6], yet the protocols governing them address connectivity and interoperability but not resource governance—how much an agent may consume or how long it may operate.
Agent Contracts address this gap by extending the contract metaphor from task allocation to resource governance. Where the Contract Net Protocol [51] asks “who should do this task?”, Agent Contracts ask “within what bounds may this task be performed?” We draw on contract theory from economics, coordination theory from distributed systems, and resource-bounded computation from real-time systems to form a formal framework for resource-bounded autonomous AI.
This paper makes two contributions. First, we define an Agent Contract as a formal tuple $C=(I,O,S,R,T,\Phi,\Psi)$ that unifies input/output specifications, resource constraints, temporal boundaries, and success criteria into a coherent governance mechanism. Second, we establish conservation laws for multi-agent systems that ensure budget discipline across delegation hierarchies, enabling composable coordination patterns where contracting itself becomes an agent capability. We validate these contributions across four experiments demonstrating 90% token reduction in iterative workflows, zero conservation violations in multi-agent delegation, and measurable quality-resource tradeoffs. Together, these contributions provide formal foundations for explicit resource governance in autonomous AI systems.
2 Theoretical Foundations
Agent Contracts draw on three theoretical traditions: contract theory from economics, multi-agent coordination from computer science, and resource-bounded computation from real-time systems.
Contract Theory. Bolton and Dewatripont (2005) [8] formalize how agreements are constructed under asymmetric information. Three concepts apply to agent governance: moral hazard (hidden actions; in LLM agents, unpredictable resource consumption), incomplete contracts (separating success criteria from execution strategies) [21], and mechanism design (specifications that elicit desired behavior). Recent work [45, 60] identifies principal-agent dynamics in LLM systems but provides conceptual rather than operational frameworks.
Multi-Agent Coordination. Classical MAS research [57, 48] established foundations for agent coordination. The Contract Net Protocol [51] demonstrated that explicit contracting could coordinate distributed systems. Coordination theory [37] identifies fundamental problems (managing shared resources, producer-consumer relationships, and simultaneity constraints), each corresponding to challenges in multi-agent LLM systems. Research on normative multi-agent systems [7] formalizes how norms govern agent behavior through obligations, prohibitions, and permissions. Agent Contracts operationalize this perspective: resource constraints function as prohibitions (agents must not exceed budgets), success criteria as obligations (agents must achieve quality thresholds), and the contract lifecycle provides regimented enforcement where violations trigger automatic termination.
A key insight from MAS theory is that coordination mechanisms must respect conservation laws—resources allocated to subtasks cannot exceed parent resources. This principle requires explicit enforcement in LLM systems where token consumption is stochastic and observable only after the fact.
Resource-Bounded Computation. Simon’s theory of bounded rationality [49, 50] established that agents with limited cognitive resources must satisfice rather than maximize. Agent Contracts operationalize satisficing by defining acceptable quality thresholds within resource budgets.
The algorithmic foundation comes from contract algorithms [63], which specify computation budgets before activation. Unlike anytime algorithms [64] that can be interrupted arbitrarily, contract algorithms enable strategic resource allocation. An Agent Contract transforms an LLM agent into a contract algorithm where bounds $R$ are known in advance. Real-time systems theory [9] contributes the distinction between hard constraints (violation causes termination) and soft constraints (permitting graceful degradation).
3 Related Work
3.1 Agent Architectures and Coordination Protocols
The development of LLM-based agents has accelerated rapidly since 2022. ReAct [59] introduced the paradigm of synergizing reasoning and acting. Chain-of-Thought prompting [56] established that computational depth correlates with output quality, motivating explicit resource governance. Toolformer [47] showed that language models can learn to use external tools, expanding the resource consumption profile of agents. Fully autonomous systems like AutoGPT [19] and Generative Agents [44] revealed both the potential and governance challenges of unbounded execution.
Coordination protocols have evolved to address different concerns. The Contract Net Protocol [51] established task allocation through bidding; MCP [4] standardizes tool connectivity between models and external resources; A2A [16] enables discovery and interoperability across heterogeneous agent systems. The recent formation of the Agentic AI Foundation [33] under the Linux Foundation—with contributions including MCP, OpenAI’s AGENTS.md, and Block’s goose—signals industry consensus on connectivity and documentation standards. However, resource governance remains outside this scope; none of these initiatives formalize how much agents may consume.
3.2 Budget-Aware Reasoning and Resource Management
A growing body of work addresses resource efficiency in LLM reasoning. The TALE framework [11] introduces token-budget-aware reasoning, achieving 68% reduction in token usage with less than 5% accuracy degradation. Critically, the authors identify “token elasticity”: LLMs often exceed specified budgets when constraints are tight, demonstrating that prompting alone is insufficient for strict enforcement. BudgetThinker [35] addresses this through control tokens injected during inference, coupled with reinforcement learning to achieve precise budget adherence. SelfBudgeter [55] enables models to predict required token budgets based on task complexity. Liu et al. (2025) [34] demonstrate that simply granting larger tool-call budgets fails to improve agent performance; their Budget-Aware Tool Selection (BATS) framework shows that explicit budget awareness enables effective scaling.
Snell et al. (2024) [52] provide a budget-aware evaluation framework, demonstrating that when compute is equalized, sophisticated reasoning strategies often do not outperform simpler baselines; much apparent improvement comes from using more resources rather than using them more intelligently. This finding underscores the importance of explicit resource accounting.
Infrastructure-level resource management has matured separately. LLM serving systems [61, 27] optimize throughput and memory efficiency at the inference layer. LLMOps platforms provide budget tracking, alerting, and rate limiting at the organizational level. However, these operate below or above the application layer; neither provides formal contracts that govern individual agent behavior within multi-agent workflows.
3.3 Agent Safety and Formal Verification
Foundational AI safety work [2, 46] identifies problems including reward hacking and safe exploration. Recent work calls for responsible LLM-empowered multi-agent systems [23], recognizing that uncertainties compound across agent interactions.
Formal verification approaches have begun addressing agentic AI specifically. Zhang et al. (2025) [62] propose a modeling framework with 17 properties for host agents and 14 for task lifecycles, expressed in temporal logic. This work is complementary to Agent Contracts. Formal verification addresses whether systems satisfy properties; Agent Contracts specify what resource constraints systems must satisfy.
3.4 Multi-Agent Coordination Frameworks
LLM-based multi-agent systems have proliferated with frameworks addressing different coordination paradigms. MetaGPT [22] integrates human workflow patterns into multi-agent collaboration, using standardized operating procedures to structure agent interactions. AutoGen [58] frames coordination as asynchronous conversation among specialized agents, with each agent capable of responding, reflecting, or invoking tools based on message content. LangGraph [28] provides graph-based orchestration with explicit state management and checkpointing, enabling durable execution of complex workflows. CrewAI [12] takes a role-based approach where agents are assigned organizational roles (researcher, developer, etc.) with corresponding capabilities.
Recent surveys characterize the landscape systematically. Tran et al. (2025) [53] analyze collaboration mechanisms across five dimensions: actors involved, interaction types (cooperative, competitive, or coopetitive), organizational structures, coordination strategies, and communication protocols. Additional surveys examine architectural patterns [38, 31] and how message-passing architectures affect coordination effectiveness [54]. These taxonomies reveal sophisticated pattern vocabulary but consistently note that resource governance remains underdeveloped. Research on self-resource allocation [32] demonstrates that LLMs can serve as effective resource allocators, with planner-based approaches outperforming real-time orchestration for concurrent task management. However, this work studies allocation capability rather than providing a formal governance framework.
To clarify this governance gap, Table LABEL:tbl-framework-comparison summarizes resource governance features across eight major agent frameworks. All provide operational controls (iteration limits, timeouts, and rate limiting), reflecting engineering best practices for preventing runaway execution. However, none provide the formal governance layer that Agent Contracts introduce: cost budgets, temporal deadlines, success criteria, or conservation laws for multi-agent delegation.
Table 1: Governance features across agent frameworks. Y = native, P = partial, – = none. Italic: operational controls. Bold: governance features unique to Agent Contracts. LG = LangGraph, AG = AutoGen, Crew = CrewAI, OAI = OpenAI Agents SDK, ADK = Google ADK, BR = Amazon Bedrock, LI = LlamaIndex, smol = smolagents. aAutoGen is being merged with Semantic Kernel into Microsoft Agent Framework. [30, 39, 13, 43, 17, 1, 36, 24]
LG AGa Crew OAI ADK BR LI smol
Max iterations Y Y Y Y Y Y Y Y
Timeout Y Y Y Y Y Y Y Y
Rate limiting Y P Y P P Y P Y
Token limits P Y Y Y Y Y Y Y
Observability Y Y Y Y Y Y Y Y
Guardrails P P Y Y Y P
Agent Contract
The table reveals a consistent pattern: existing frameworks provide operational controls but lack formal governance mechanisms. Recent practitioner perspectives frame “agent engineering” as iterative refinement for reliability [29], but do not address resource governance. The following section presents Agent Contracts as a framework that fills this gap.
4 The Agent Contract Framework
4.1 Contract Definition
An Agent Contract $C$ is defined as a seven-tuple:
$C=(I,O,S,R,T,\Phi,\Psi)$
The components capture the complete specification for bounded agent execution. The input specification $I$ defines the schema and constraints for acceptable inputs. The output specification $O$ defines the schema and quality criteria for deliverables. The skill set $S$ enumerates the capabilities (tools, functions, and knowledge domains) available to the agent. Resource constraints $R$ specify a multi-dimensional budget governing consumption. Temporal constraints $T$ establish time-related boundaries and duration limits. Success criteria $\Phi$ define measurable conditions for contract fulfillment. Finally, termination conditions $\Psi$ specify events that end the contract regardless of fulfillment.
This formulation synthesizes contract theory in economics, where contracts align incentives and define obligations between parties [8], with real-time systems theory, where correctness depends on meeting explicit timing constraints [9]. The contract serves as both specification (defining what the agent should do) and governance mechanism (constraining how the agent may operate).
An important distinction separates $I$ from $R$. The input specification $I$ defines what the agent receives: task content, context, and parameters. The resource constraints $R$ define how much the agent may consume while processing: token budgets, API call limits, time bounds. An agent may receive a small input but consume many resources through complex reasoning, or receive a large input but consume few resources through simple transformation.
4.2 Contract Components
Input and Output Specifications. The input specification $I=(\sigma_{I},\mathcal{V}_{I},\mathcal{P}_{I})$ comprises the input schema $\sigma_{I}$, validation rules $\mathcal{V}_{I}$, and preprocessing transformations $\mathcal{P}_{I}$. For example, a code review agent might specify $\sigma_{I}$ as {repository: string, pr_id: integer}, $\mathcal{V}_{I}$ as pr_id > 0, and $\mathcal{P}_{I}$ as fetch_diff(pr_id).
The output specification $O=(\sigma_{O},Q_{min},\mathcal{F}_{O})$ comprises the output schema $\sigma_{O}$, minimum acceptable quality threshold $Q_{min}$, and formatting requirements $\mathcal{F}_{O}$. For the code review agent, $\sigma_{O}$ might be {summary: string, issues: list, approval: boolean}, $Q_{min}=0.8$ (requiring 80% of issues correctly identified), and $\mathcal{F}_{O}$ as markdown format with severity labels.
Skills and Capabilities. The skill set $S=\{s_{1},s_{2},...,s_{m}\}$ where $s_{i}\in\mathcal{S}_{available}$ enumerates what the agent can do [5]. Each skill $s_{i}$ may have associated costs $c(s_{i})$ and success probabilities $p(s_{i})$. Skills encompass tool invocations (web search, code execution, API calls), knowledge domains (legal, medical, technical), and cognitive capabilities (reasoning, planning, summarization). The contract restricts the agent to skills in $S$; for example, an agent contracted for data analysis cannot invoke payment APIs even if technically accessible.
Resource Constraints. The resource constraint $R=\{r_{1},r_{2},...,r_{n}\}$ defines a multi-dimensional budget. Common resource dimensions include:
Resource Symbol Unit Example
LLM Tokens $r_{tok}$ tokens 100,000
API Calls $r_{api}$ calls 50
Iterations $r_{iter}$ rounds 10
Web Searches $r_{web}$ queries 10
Compute Time $r_{cpu}$ seconds 300
External Cost $r_{cost}$ USD 5.00
Each resource $r_{i}$ has a budget $b_{i}$ and consumption function $c_{i}(t)$. Constraint satisfaction requires $\forall i:c_{i}(t)\leq b_{i}$.
Temporal Constraints. The temporal constraint $T=(t_{start},\tau)$ comprises the contract activation timestamp $t_{start}$ and the contract duration (time-to-live) $\tau$. The constraint requires $t_{current}-t_{start}\leq\tau$. While users often think in terms of deadlines (“complete by 5pm”), duration is the operational primitive—a deadline is simply $t_{deadline}=t_{start}+\tau$.
Success Criteria and Termination. Success criteria $\Phi=\{(\phi_{1},w_{1}),(\phi_{2},w_{2}),...,(\phi_{k},w_{k})\}$ pair measurable conditions $\phi_{i}$ with weights $w_{i}$. The contract is fulfilled when $\sum w_{i}\cdot\mathbb{1}[\phi_{i}]\geq\theta$ for threshold $\theta$. Conditions may include task completion (all_items_processed), quality metrics (accuracy > 0.95), or business logic (response_generated AND reviewed).
The output quality threshold $Q_{min}$ (in specification $O$) is a structural requirement on output format and minimum acceptability. Success criteria $\Phi$ are fulfillment conditions that may include quality checks (e.g., $\phi_{1}=Q\geq Q_{min}$) along with other conditions. An agent might produce output meeting $Q_{min}$ but fail $\Phi$ if other criteria are unmet.
Termination conditions $\Psi=\{\psi_{1}\lor\psi_{2}\lor...\lor\psi_{l}\}$ define when the contract ends regardless of success. Common termination conditions include resource exhaustion ($\exists r_{i}:c_{i}\geq b_{i}$), duration expiration ($t-t_{start}>\tau$), explicit cancellation (external signal), and unrecoverable errors (critical failure states). The existential quantifier ensures that exceeding any resource budget causes termination, not just aggregate overruns.
4.3 Contract Lifecycle
Contracts progress through distinct states following the transition pattern DRAFTED $\rightarrow$ ACTIVE $\rightarrow$ {FULFILLED, VIOLATED, EXPIRED, TERMINATED}. A contract begins in the DRAFTED state, where its parameters are specified but execution has not begun. Activation transitions the contract to ACTIVE, at which point resources are reserved and monitoring begins. From ACTIVE, the contract reaches exactly one of four terminal states.
The terminal state FULFILLED indicates that success criteria $\Phi$ were satisfied within all resource and temporal constraints. VIOLATED indicates that some constraint in $R$ or $T$ was breached before success criteria were met. EXPIRED indicates that the duration $\tau$ was exceeded. TERMINATED indicates external cancellation, regardless of progress toward success criteria.
Transitions between states are governed by formal guard conditions:
From To Guard Condition
DRAFTED ACTIVE $\text{activate}()\land\text{resources\_available}()$
ACTIVE FULFILLED $\sum w_{i}\cdot\mathbb{1}[\phi_{i}]\geq\theta$
ACTIVE VIOLATED $\exists r_{i}:c_{i}\geq b_{i}$
ACTIVE EXPIRED $t-t_{start}>\tau$
ACTIVE TERMINATED $\text{cancel\_signal}()$
The lifecycle model ensures clear accountability. Every contract reaches exactly one terminal state, enabling unambiguous resource release and audit logging.
5 Resource Tracking and Monitoring
The preceding chapter defined what contracts are: their structure, components, and lifecycle. This chapter addresses how resources are tracked during execution. While the framework tracks multiple resource types (tokens, API calls, tool invocations, compute time, cost), we focus here on two foundational aspects: how token budgets decompose into measurable categories, and how runtime monitoring provides visibility into constraint utilization.
5.1 Token Budget Decomposition
Modern LLMs distinguish input, reasoning, and output tokens. We model this as $R_{tok}=(r_{in},r_{r},r_{out})$. Since input tokens $r_{in}$ are largely determined by task context, the controllable budget is $B_{ctrl}=B_{tok}-r_{in}$, representing tokens available for reasoning and output. This decomposition enables fine-grained monitoring across categories, supporting both real-time adaptation and post-hoc analysis.
5.2 Runtime Monitoring
During execution, a monitoring system tracks both resource consumption and temporal progress in real-time. The monitor function $\text{Monitor}:(C,t)\rightarrow(\vec{c},\vec{u},\tau_{util})$ takes a contract $C$ and current time $t$ and returns three values: the resource consumption vector $\vec{c}$, the resource utilization vector $\vec{u}=\vec{c}/\vec{b}$ (computed element-wise), and the duration utilization $\tau_{util}=(t-t_{start})/\tau$ ranging from 0 to 1. The agent can query these values at any time to adapt its strategy as constraints tighten. Note that utilization is monotonically non-decreasing since resource consumption is cumulative.
A useful aggregate metric captures the most-constrained resource:
$\text{utilization}(t)=\max\left(\frac{t_{current}-t_{start}}{\tau},\max_{i}\frac{c_{i}(t)}{b_{i}}\right)$
This single value summarizes how close the agent is to any constraint boundary, enabling simple threshold-based policies (e.g., “warn when utilization exceeds 80%”) without requiring sophisticated optimization.
Communicating Budget to Agents. A current approach is budget-aware prompting: injecting remaining budget into system prompts or providing dynamic status updates during execution. This enables agents to self-regulate—producing concise outputs when budget is tight or taking exploratory actions when resources are ample. As the field matures, native solutions may emerge (e.g., models with built-in resource awareness), but prompt-based communication provides a practical mechanism with current infrastructure.
6 Multi-Agent Coordination Under Contracts
The preceding sections establish contracts for individual agents. However, complex tasks often require multiple agents working together: a researcher gathering data, an analyzer identifying patterns, a reporter synthesizing findings. This extension from single-agent to multi-agent governance raises new questions: How should a parent budget be divided among child agents? What happens when one agent exceeds its allocation while others remain under budget? How can the system guarantee that aggregate consumption respects the original constraint?
These questions have theoretical grounding in coordination theory. Malone and Crowston (1994) [37] identified managing shared resources as a fundamental coordination problem. In LLM-based multi-agent systems, the shared resource is typically the token budget (and associated cost), but the same principles apply to API call limits, compute time, and other constrained resources. The key insight is that contracts provide a natural unit of delegation. When an orchestrator creates subcontracts for workers, the contract specification ensures that each worker operates within defined bounds, and the aggregate of those bounds respects the parent constraint.
6.1 Conservation Laws and Budget Allocation
When multiple agents collaborate, contracts must govern how resources flow between them. The fundamental constraint is conservation: total consumption cannot exceed the system budget:
$\sum_{j\in\text{agents}}c_{j}^{(r)}\leq B^{(r)}\quad\forall r\in R$
This invariant holds regardless of execution pattern—sequential, parallel, hierarchical, or competitive.
Initial Allocation. Before execution, the total budget $B$ must be divided among agents. Three strategies apply depending on available information:
  • Proportional: $b_{j}=\omega_{j}/\sum\omega_{k}\cdot(B-B_{reserve})$, where $\omega_{j}$ reflects estimated task complexity
  • Equal: $b_{j}=(B-B_{reserve})/n$, when complexity is unknown
  • Negotiated: Agents request budgets; a coordinator allocates based on requests with caps to prevent over-claiming
A reserve buffer $B_{reserve}$ (typically 10–15%) accommodates coordination overhead and unexpected costs.
Dynamic Reallocation. As agents complete, unused budget returns to a shared pool:
$B_{available}(t)=B_{reserve}+\sum_{j\in\text{completed}}(b_{j}-c_{j})$
This enables budget pooling—efficient agents effectively subsidize those requiring more resources, improving overall throughput while maintaining total budget discipline.
6.2 Coordination Patterns Through a Contract Lens
Recent work has identified recurring design patterns for agentic AI systems, including routing, orchestration, parallelization, and iterative refinement [41, 3, 18].
We organize these patterns through a contract-centric lens, focusing on how resource constraints govern each pattern’s behavior and where contracts provide critical safety guarantees. We focus here on two typical control flow patterns, task routing and delegation decisions, since these are where contracts add the most value. Execution within any pattern can be sequential or parallel; these are orthogonal concerns that compose naturally (as we discuss below).
Routing. An input classifier directs requests to specialized handlers based on task characteristics, selecting the best-suited agent for each input and allocating the corresponding budget [3]. Budget is reserved per potential branch; unused allocations return to the pool. This enables separation of concerns— specialized agents outperform generalists on their specific tasks.
When specialists have explicit contracts, routing becomes more principled: the router matches task requirements against specialist capabilities, resource profiles, and success criteria. Furthermore, with well-defined contracts, the router need not be limited to a fixed pool; it can dynamically instantiate or configure an agent specifically for the required contract. This blurs the line between routing and orchestration, with the contract serving as the specification for agent creation, not just agent selection.
Orchestrator-Workers. A central orchestrator dynamically decomposes tasks and delegates to worker agents, synthesizing results [3, 18]. This pattern can extend to multiple levels (hierarchical orchestration).
From a contract perspective, this pattern is particularly significant: the orchestrator drafts and issues subcontracts to workers. Each subcontract specifies the worker’s task, allocated budget, and success criteria:
$\text{orchestrator}(C_{parent})\rightarrow\{C_{i}=(I_{i},O_{i},S_{i},R_{i},T_{i},\Phi_{i},\Psi_{i})\}_{i=1}^{k}$
This frames contracting as a capability: the orchestrator must understand the contract framework to effectively delegate work. The conservation law (Section 6.1) constrains subcontract allocation: $\sum R_{i}\leq R_{parent}$.
The implications of contracting as a capability are significant:
Implication Description
Recursive delegation Agents spawn sub-agents that themselves have contracting capability, enabling hierarchical self-organization
Bounded autonomy Even highly capable orchestrators remain governed by their parent contract: they can create subcontracts but cannot exceed their own constraints
Dynamic team formation Agents form coalitions and delegate work without centralized coordination, as long as conservation laws are satisfied
Dynamic agent instantiation Rather than selecting from a fixed pool, agents instantiate specialists on-demand; the contract becomes the specification for agent creation, not just selection
Meta-governance Contracts can govern the creation of other contracts, enabling principled scaling of multi-agent systems
These patterns demonstrate how contracts can govern increasingly complex multi-agent systems. However, practical enforcement faces fundamental constraints that shape what contracts can and cannot guarantee.
7 Fundamental Limitations and Practical Enforcement
7.1 Single-Call Enforcement Constraints
A critical limitation exists. Token consumption is only known after an LLM call completes, not during execution:
$c_{tok}(t)=\begin{cases}\text{unknown}&\text{during API call}\\ c_{actual}&\text{after API returns}\end{cases}$
Even streaming APIs provide total usage metadata only after generation completes. This asymmetry has three important consequences: contracts cannot prevent a single expensive call from exceeding budget, contracts can prevent subsequent calls after budget is exceeded, and the primary value is therefore multi-call protection.
7.2 Enforcement Capabilities
Despite single-call limitations, contracts provide substantial value across multiple enforcement dimensions:
Constraint Enf. Mechanism
Multi-call budgets Full Stop after cumulative threshold
Iteration limits Full Count and halt at $r_{iter}$
API call limits Full Count external invocations
Duration limits Full Check elapsed time before each action
Cost ceilings Approx Track cumulative cost; bound via max_tokens
Contracts provide particularly high value in five scenarios. First, retry loops benefit from contracts that prevent runaway costs when tasks fail repeatedly. Second, validation cycles benefit from contracts that limit iterative refinement (e.g., code generation $\rightarrow$ test $\rightarrow$ fix loops). Third, multi-agent workflows benefit from contracts that stop downstream agents when upstream exceeds budget. Fourth, tool-heavy agents benefit from contracts that control cumulative cost of web searches and API calls. Fifth, long-running sessions benefit from contracts that enforce session-level budgets across many interactions.
7.3 Future Infrastructure Requirements
True real-time enforcement would require further API-level changes. While some providers now offer cumulative token counts during streaming, key capabilities remain missing: interruptible generation allowing mid-generation cancellation with partial output, token reservation to pre-allocate budget with guaranteed hard limits, and budget-aware inference enabling models to respect token budgets natively, as explored in recent work on budget-aware reasoning [11, 35].
Such infrastructure would enable contracts to provide hard guarantees rather than best-effort enforcement. Until then, contracts remain most valuable for multi-call and multi-agent scenarios.
8 Empirical Evaluation
We validate Agent Contracts through four complementary experiments spanning single-agent and multi-agent settings. Each experiment targets specific framework components; collectively they validate the enforcement mechanisms ($R$, $T$, $\Phi$, $\Psi$) and coordination primitives (conservation laws, contract delegation) that distinguish Agent Contracts from existing approaches. Experiments use a reference implementation with Google ADK (Code Review, Research Pipeline, Crisis Communication) or LiteLLM (Strategy Modes), using Gemini 2.5 Flash and Flash-Lite (knowledge cutoff: January 2025). [^1] Implementation available at https://github.com/flyersworder/agent-contracts. The framework is under active development; we provide experiment code and data for reproducibility. Statistical analysis employs bootstrap confidence intervals (10,000 resamples) with BCa correction.
8.1 Experimental Overview
The following table summarizes our experimental design, with each experiment validating a distinct aspect of the framework.
Experiment n Validates Key Result
Code Review 70 Runaway prevention 90% token reduction
Research Pipeline 50 Conservation laws 0 violations
[…]
Crisis Comm. 24 Failure prevention 23% fewer tokens
8.2 Runaway Prevention in Iterative Workflows
[…]
Design. Each problem is executed twice in a within-subjects design: CONTRACTED (50K token budget, max 3 iterations) versus UNCONTRACTED (no token limits, max 6 iterations). CONTRACTED agents receive budget-aware prompts and dynamic status updates showing both token consumption and iteration progress; UNCONTRACTED agents use identical base prompts without resource awareness.
[…]
Metric UNCONTRACTED CONTRACTED Change $p$-value
Token Usage 34,606 3,461 -90% 0.0007***
[…]
Iterations 3.00 1.71 -43% <0.0001***
LLM Calls 9.0 4.5 -50% <0.0001***
Success Rate 60.0% 52.9% -7.1pp 0.13 (NS)
A complementary single-agent experiment with 24 time-critical crisis communication scenarios provides additional evidence: agents with explicit quality thresholds ($Q\geq 0.80$) and iteration limits achieved 23% token reduction ($p=0.005$) with statistically equivalent quality ($p=0.32$). Notably, one UNCONTRACTED agent failed entirely—stuck in an evaluation loop without submitting output—while the CONTRACTED version succeeded, demonstrating that iteration governance prevents agent failures, not just improves efficiency.
8.3 Conservation Laws in Multi-Agent Coordination
We evaluate conservation laws and contract delegation (Section 6) using a three-agent research pipeline: Researcher $\rightarrow$ Analyzer $\rightarrow$ Reporter. The orchestrator delegates sub-contracts to each worker via DelegatingAdkAgent, enforcing both the conservation invariant $\sum b_{i}\leq B$ and per-tool limits at allocation time.
Design. Fifty research topics across five categories (technology, science, business, health, society) are evaluated. CONTRACTED agents receive explicit budget allocations with per-tool limits (e.g., Researcher limited to 6 web searches) and budget-aware prompts; UNCONTRACTED agents operate without constraints. Quality is assessed via multi-judge LLM evaluation following best practices for rating indeterminacy [20].
[…]
8.4 Quality-Resource Tradeoffs via Contract Modes
Contract modes operationalize Simon’s satisficing principle [49]: agents achieve acceptable quality within defined resource bounds. We test whether different contract configurations produce measurable behavioral differences using logic reasoning problems from OpenR1 (released February 2025, after model cutoff).
Design. Fifty medium-difficulty logic puzzles are evaluated under three contract modes: URGENT (no extended reasoning, 30s timeout), ECONOMICAL (low reasoning effort, 60s timeout), and BALANCED (medium reasoning effort, 90s timeout). Success requires exact numeric answer match.
Results. Contract modes produce a clear quality-resource gradient. BALANCED mode achieves 86% success rate versus 70% for URGENT, investing 75% more tokens for 16 percentage points higher success ($p\approx 0.05$). The reasoning_effort parameter provides direct control: users requiring speed accept lower accuracy (URGENT); those requiring accuracy invest more resources (BALANCED).
Mode Success Rate Reasoning Tokens Avg Time Timeout Rate
URGENT 70% 0 6.9s 26%
ECONOMICAL 76% 718 12.5s 14%
BALANCED 86% 1,519 16.9s 10%
[…]
Framework Component Experiment Validation
Budget awareness All experiments Total budget in prompts; dynamic status updates (Code Review)
Resource constraints $R$ Code Review Tokens, iterations, LLM calls
Agent delegation Research Pipeline Sub-contracts, conservation laws, per-tool limits
Success criteria $\Phi$ Crisis Comm. Quality thresholds
Contract modes Strategy Modes URGENT/ECONOMICAL/BALANCED
9 Conclusion
Agent Contracts provide a formal framework for governing autonomous AI agents through explicit resource and temporal constraints. The contract specification $C=(I,O,S,R,T,\Phi,\Psi)$ unifies resource, temporal, and quality governance with conservation laws for multi-agent delegation.
[…]
Several extensions merit investigation: runtime cancellation would require API-level support from model providers to halt mid-execution; learning-based contract design would enable agents to predict budgets and draft subcontracts [11, 35]; human-in-the-loop contracts would specify when approval is required; and milestone-based governance would support continuous agent operation. As agentic AI moves to production, formal governance becomes essential—not only for cost control but for accountability and value alignment, ensuring agents operate within bounds acceptable to stakeholders. Agent Contracts provide one such foundation.